Quantitative Problem: You need $14,000 to purchase a used car. Your wealthy uncle is willing to lend you the money as an amortized loan. He would like you to make annual payments for 4 years, with the first payment to be made one year from today. He requires a 5% annual return. What will be your annual loan payments? Round your answer to the nearest cent. Do not round intermediate calculations. $ How much of your first payment will be applied to interest and to principal repayment? Round your answer to the nearest cent. Do not round intermediate calculations. Interest: $ Principal repayment: $
Answer #1

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

14000=Annuity[1-(1.05)^-4]/0.05

14000=Annuity*3.545950504

Annuity=14000/3.545950504

=$3948.17(Approx)=annual loan payments.

Interest payment=$14000*5%=$700

Principal repayment=(3948.17-700)=$3248.17(Approx).